OpenStudy (anonymous):

What is the value of y in the following system of equations? 2x+3y=6 2x+y=-2 1)1 2)2 3)-3 4)4

OpenStudy (saifoo.khan):

-3

OpenStudy (amistre64):

-2y = 10 y = -5 ... but that aint an option i sit lol 2x +3y =6 -2x -y = 4 ----------- 2y = 10 y = 5

OpenStudy (amistre64):

ack.. i forgot how to multiply

OpenStudy (amistre64):

4

OpenStudy (amistre64):

6+2 = 8 2y = 8 y = 4

OpenStudy (toxicsugar22):

y=4